OpenStudy (anonymous):

Solve the equation; check for extraneous solutions. 6|6 - 4x| = 8x + 4 Please help, this is so confusing!

OpenStudy (goformit100):

First of all @peyxo : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

\[\left| 6-4x \right| =6-4x ,if 6-4x >0 ,or 6> 4x,or4x<6,x<\frac{ 6 }{4 },x<\frac{ 3 }{2 }\] \[solve 6(6-4x)=8x+4, but keep inmind x <\frac{ 3 }{ 2 }\] \[similarly \left| 6-4x \right|=-\left( 6-4x \right), if 6-4x<0,6<4x,4x>6, x>\frac{ 6 }{4 },x>\frac{ 3 }{ 2 }\] \[solve -6(6-4x)=8x+4,but keep inmind x >\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

if it is absolute value you make it posotive in the inside lines so it would be 7 the answer cause after that you jusgt simplify:) hope I helped